THE OBSERVATION AND PROPERTIES OF (H_{2}O⋯HCl)⋯Ar

\underline{Z. KISIEL}, L. PSZCZÓŁKOWSKI, E. BIAŁKOWSKA-JAWORSKA, O. DESYATNYK, M. J. NOWAK, Institute of Physics, Polish Academy of Sciences, Al. Lotników 32/46, 02-668 Warszawa, Poland.

The trimer (H_{2}O⋯HCl)⋯Ar was observed in this laboratory by supersonic expansion cavity-FTMW spectroscopy, during the work carried out on H_{2}O...HCl^{a} and (H_{2}O)_{2}HCl.^{b} The title trimer appears to be useful test-bed for detailed assessment of the effect of complexing a Rare-gas atom to the more strongly bound H_{2}O⋯HCl subunit. (H_{2}O⋯HCl)⋯Ar is nearplanar and since the H_{2}O⋯HCl unit is oriented at ≈ 27^{∘} to the b-inertial axis both the μ_{b} and the μ_{a} dipole moment components are appreciable, giving rise to a rich rotational spectrum. Presently we present a summary of the information obtained for this trimer. Five different isotopomers of the trimer, resulting from ^{37}Cl, D, and ^{18}O substitution, have been observed. In addition to the usual spectroscopic constants it was possible to determine accurate values of the χ_{ab} nuclear quadrupolar splitting constant for the chlorine nucleus, and thus the principal nuclear quadrupole tensor and the associated rotation angle. Electric dipole moment of the trimer was measured and supporting ab initio calculations at the MP2/aug-cc-pVTZ level were carried out. Following a recent test of the applicability of the r_{m} method to weakly bound complexes^{c} we derived such a geometry for (H_{2}O⋯HCl)⋯Ar and find it to be in satisfactory agreement with the ab inito calculations. Comparison of results obtained for the trimer with the accurate data available for H_{2}O⋯HCl allow the effects of this first step in inert-gas matrix isolation of H_{2}O⋯HCl to be identified.

aZ. Kisiel, B. A. Pietrewicz, P. W. Fowler, A. C. Legon, E. Steiner, J. Phys. Chem. A 104, 6970 (2000)

bZ. Kisiel, E. Białkowska-Jaworska, L. Pszczółkowski, A.Milet, C.Struniewicz, R.Moszynski, J.Sadlej, J. Chem. Phys. 112, 5767 (2000)

cZ. Kisiel, J. Mol. Spectrosc. 218, 58 (2003)